OpenStudy (anonymous):

how do I solve COT pi/4 without a calculator

OpenStudy (anonymous):

special angles can be memorized

OpenStudy (anonymous):

you are looking for \(\frac{\cos(\frac{\pi}{4})}{\sin(\frac{\pi}{4})}\)

OpenStudy (anonymous):

if you know what those values are, you win, otherwise you need to look them up

OpenStudy (anonymous):

is 1 right..